Pam is half as old as her brother Bill. Six year ago Bill was four times older than Pam. How old is each now?
Here’s what we did:
2P=B
B-6=4P
But, in solving this, we get
B=4P+6
2P=4P+6
-2P=6, P=-3 which isn’t right
Where are we going wrong?

Answer by Earlsdon(6294)

Your first line is ok!
B = 2P
In your second line you have neglected the fact that 6 years ago, both Pam and Bill were 6 years younger, so your second line should read:
B-6 = 4(P-6) Now substitute for B, B = 2P and expand the right side..
2P-6 = 4P-24 Subtract 2P from both sides.
-6 = 2P-24 Add 24 to both sides.
18 = 2P Finally, divide both sides by 2.
P = 9 and...
B = 2P
B = 19